OpenStudy (cedric):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

OpenStudy (cedric):

i wana do the last two

OpenStudy (anonymous):

Go ahead then...

hero (hero):

What's the matter? Do you know how to factor them or not?

OpenStudy (cedric):

No, and nice comment @waterineyes

OpenStudy (anonymous):

Thank you @cedric

OpenStudy (anonymous):

What you know about factorization ???

OpenStudy (cedric):

ok, let e try real quick for the last one

OpenStudy (anonymous):

Take your time..

OpenStudy (cedric):

8x2 − 11x + 3 also equals x(8x - 11) + 3 ?

OpenStudy (cedric):

then you'd find the gcf

OpenStudy (cedric):

and i dont think there is one, so you'd factor the 3 right?

OpenStudy (cedric):

am i good so far?

OpenStudy (cedric):

well then i think you'd end up with (x - 1)(8x - 3)

OpenStudy (anonymous):

Wow...

OpenStudy (cedric):

Yo look i dont have time for you to act smart, i jusrt need help

OpenStudy (anonymous):

Do for second last now..

OpenStudy (anonymous):

Don't you have faith on yourself?/ you are right in doing that..

OpenStudy (anonymous):

Your answer is right but I did not get your way how you solved it..

hero (hero):

Cedric, explain how you got from x(8x - 11) + 3 to (x - 1)(8x - 3) again.

OpenStudy (cedric):

ok imma take a pic of my paper, im doiing it on

OpenStudy (anonymous):

Hero is like my Elder Brother so follow him properly..

OpenStudy (cedric):

8x2 - 11x + 3 = 8(x - 1)(x - 3/8) then multply by 8

hero (hero):

Cedric, I appreciate you posting that, but you need to explain your steps a little clearer. I'm having trouble following them. How did you get from 8x^2 - 11x + 3 to 8(x - 1)(x - 3/8) ?

OpenStudy (cedric):

by simplifying it

OpenStudy (cedric):

oh shiz my bad use this formula ax2 + bx + c = a(x - x1)(x - x2)

hero (hero):

I'm not really following you. Can you explain it to me on vyew?

OpenStudy (cedric):

vyew?

hero (hero):

Yes, it is an online collaboration site.

OpenStudy (cedric):

seems legit

hero (hero):

Do you have a headset with a microphone?

hero (hero):

Cedric?